# Connections between Combinations Without Specified Separations and Strongly Restricted Permutations, Compositions, and Bit Strings

### v2

@@ Abstract
Let $S_{n}$ and $S_{n,k}$ be, respectively, the number of subsets and $k$-subsets of $\mathbb{N}_{n}=\{1,\ldots,n\}$ such that no two subset elements differ by an element of the set $\mathcal{Q}$, the largest element of which is $q$. We prove a bijection between such $k$-subsets when $\mathcal{Q}=\{m,2m,\ldots,jm\}$ with $j,m>0$ and permutations $\pi$ of $\mathbb{N}_{n+jm}$ with $k$ excedances satisfying $\pi(i)-i\in\{-m,0,jm\}$ for all $i\in\mathbb{N}_{n+jm}$. We also identify a bijection between another class of restricted permutation and the cases $\mathcal{Q}=\{1,q\}$. This bijection allows us to prove a conjectured recursion relation for the number of such permutations which corresponds to the case $\mathcal{Q}=\{1,4\}$. We also obtain the generating function for $S_{n}$ in the case $\mathcal{Q}=\{1,5\}$ by first obtaining generating functions for the numbers of closed walks of a given length on a particular class of directed pseudograph. We give some classes of $\mathcal{Q}$ for which $S_{n}$ is also the number of compositions of $n+q$ into a given set of allowed parts. A bijection between the $k$-subsets for any $\mathcal{Q}$ and bit strings is also noted. Aided by this, an efficient algorithm for finding $S_{n}$ and $S_{n,k}$ is given. We also prove a bijection between $k$-subsets for a class of $\mathcal{Q}$ and the set representations of size $k$ of equivalence classes for the occurrence of a given length-($q+1$) subword within bit strings. We then formulate a straightforward procedure for obtaining the generating function for the number of such equivalence classes.
@@ 1 Introduction
When there are no restrictions, we just have the number of ordinary combinations and so $S^{\{\}}_{n}=2^{n}$ (A000079 in the OEIS [20]) and $S^{\{\}}_{n,k}=\tbinom{n}{k}$ (A007318). It appears to be well-known that $S^{\{1,\ldots,q\}}_{n}=S^{\{1,\ldots,q\}}_{n-1}+S^{\{1,\ldots,q\}}_{n-q-1}+\delta_{n+q,0}$, where $\delta_{i,j}$ is 1 if $i=j$ and 0 otherwise, and $S^{\{1,\ldots,q\}}_{n,k}=\tbinom{n+q(1-k)}{k}$. These reduce to the classic results $S^{\{1\}}_{n}=F_{n+2}$, where $F_{n}$ is the $n$-th Fibonacci number (A000045, $F_{n}=F_{n-1}+F_{n-2}+\delta_{n,1}$, $F_{n<1}=0$), and $S^{\{1\}}_{n,k}=\tbinom{n+1-k}{k}$ (A011973) [11]. Expressions for $S_{n}^{\mathcal{Q}}$ and $S^{\mathcal{Q}}_{n,k}$ have also been obtained when $\mathcal{Q}=\{q\}$ [13, 19, 15, 14, 3] and $\mathcal{Q}=\{m,2m,\ldots,jm\}$ with $j,m>0$ [17, 1].
Recently, a bijection between the $k$-subsets of $\mathbb{N}_{n}$ with disallowed differences set $\mathcal{Q}$ of which the largest element is $q$ and the restricted-overlap tilings of an $(n+q)$-board (an $(n+q)\times 1$ board) with squares and $k$ combs was found. This enabled generating functions for $S_{n}$ and recursion relations for $S_{n,k}$ to be readily obtained for a variety of families of $\mathcal{Q}$ [2]. A $(w_{1},g_{1},w_{2},\ldots,g_{t-1},w_{t})$-comb is a tile composed of $t>0$ rectangular sub-tiles (called teeth) arranged in such a way that the $i$-th tooth, which has dimensions $w_{i}\times 1$, is separated from the $(i+1)$-th by a $g_{i}\times 1$ gap [2]. The length of a comb is the sum of the lengths of the teeth and gaps. Each $\mathcal{Q}$ corresponds to a length-$(q+1)$ comb $C$ (which we will refer to as a $\mathcal{Q}$-comb) defined as follows. We label the cells of $C$ from 0 to $q$. Cell 0 is always part of the first tooth. Cell $i$ (where $i=1,\ldots,q$) is part of a tooth iff $i\in\mathcal{Q}$. Note that we can always include 0 in $\mathcal{Q}$ (and we will do so occasionally for reasons of symmetry) since subset elements are distinct and so cannot differ by 0. This 0 element can then be regarded as corresponding to cell 0 of the comb. To ensure that there is only one way to specify the $w_{i}$ and $g_{i}$ of the $\mathcal{Q}$-comb, we insist that all teeth and gaps have positive integer length. A restricted-overlap tiling using squares and $\mathcal{Q}$-combs means that any comb can be placed so that its non-leftmost cells (i.e, cells 1 to $q$) overlap the non-leftmost cells of any other combs on the board and no other forms of overlap of tiles is allowed [2] (see Fig. 6 for examples). The following theorem is a consequence of the bijection.
@@
If $B^{\mathcal{Q}}_{n}$ is the number of restricted-overlap tilings of an $n$-board with $\mathcal{Q}$-combs and squares, and $B^{\mathcal{Q}}_{n,k}$ is the number of such tilings that contain $k$ combs, then $S^{\mathcal{Q}}_{n}=B^{\mathcal{Q}}_{n+q}$ and $S^{\mathcal{Q}}_{n,k}=B^{\mathcal{Q}}_{n+q,k}$.
@@
A cycle that does not include the 0 node is called an inner cycle (e.g., $SSC_{[4]}$ in Fig. 4, where including the length increment subscript quickly identifies which comb arc we are referring to). There are a finite number of possible metatiles iff the digraph lacks an inner cycle since a walk can traverse an inner cycle an arbitrary number of times before returning to the 0 node. If all the inner cycles have at least one node in common, that node is called a common node (e.g., the $01^{q-p}$ node is the common node in the Fig. 7 digraph; either the $10^{2}$ 1 or the 01 node can be chosen as the common node in the digraph in Fig. 3; Fig. 4 and Fig. 10 show digraphs that have inner cycles but no common node). For digraphs with a common node, there are simple expressions for the recursion relations for the numbers of tilings $B^{\mathcal{Q}}_{n}$ and $B^{\mathcal{Q}}_{n,k}$ in terms of lengths of various cycles in the digraph [9, 2] (we rederive these in the form of generating functions in §2). If the digraph possesses inner cycles but no common node, as far as we are aware, there is only one other class of digraph (which has 3 inner cycles) where an analogous simple procedure has been found for obtaining the recursion relations [1].
@@
A systematic search of the OEIS for sequences $(S^{\mathcal{Q}}_{n})_{n\geq 0}$ for all $\mathcal{Q}$ with $q\leq 4$ showed that there are various connections between some types of restricted combinations and some types of restricted permutations. Relations to compositions and bit strings were also found. Here, in §3, we give bijections between between two classes of strongly restricted permutations and combinations with $\mathcal{Q}=\{1,q\}$ and $\mathcal{Q}=\{m,2m,\ldots,jm\}$. As a consequence of the former bijection, we can prove a conjectured recursion relation for the permutations corresponding to the $\mathcal{Q}=\{1,4\}$ case. In §2, we derive generating functions for $B^{\mathcal{Q}}_{n}$ and $B^{\mathcal{Q}}_{n,k}$ for tilings with metatiles generated by a particular class of digraph (which includes the 3 inner cycle class with no common node dealt with previously [1]) without a common node. This enables us to also obtain a generating function for $S^{\{1,5\}}_{n}$. In §4, although it is not all that surprising given the tiling interpretation of restricted combinations, for completeness, we also show that $S^{\mathcal{Q}}_{n}$ is the number of compositions of $n+q$ into certain sets of parts for some classes of $\mathcal{Q}$. An elementary connection between all instances of $\mathcal{Q}$ and bit strings is pointed out in §5. This forms the basis for an algorithm for efficiently calculating $S_{n}$ and $S_{n,k}$ for any given $\mathcal{Q}$ as detailed via a heavily annotated C program listed in the Appendix. In §5 we also give a bijection between $k$-subsets for certain $\mathcal{Q}$ and equivalence classes for the occurrence of certain subwords within bit strings. The bijection enables us to show that a generating function for the number of equivalence classes is always straightforward to obtain.
@@ 2 Generating functions for numbers of walks on digraphs with a pseudo-common node
A form of the following theorem was originally stated in [9], and a more compact version re-derived in [2]. Here, using the above three lemmas, we obtain the result in terms of a bivariate generating function. We will use this theorem in §3 to obtain the $S^{\{1,4\}}_{n}$ generating function.
@@
Let $g^{\mathcal{Q}}(x)$ and $g^{\mathcal{Q}}(x,y)$ denote the generating functions for $S^{\mathcal{Q}}_{n}$ and $S^{\mathcal{Q}}_{n,k}$, respectively. Then
@@ 3 Bijections between strongly restricted permutations and restricted combinations
A strongly restricted permutation $\pi$ of the set $\mathbb{N}_{n}$ is a permutation for which the number of permissible values of $\pi(i)-i$ for each $i\in\mathbb{N}_{n}$ is less than a finite number independent of $n$ [16]. Here we give two bijections between restricted combinations and types of strongly restricted permutations where the permutations satisfy $\pi(i)-i\in\mathcal{D}$. Our first bijection concerns a strongly restricted permutation where $\mathcal{D}=\{-1,0,1\}$ and any $m$ consecutive $\pi(i)$ cannot differ from one another by more than $m$.
@@
There is a bijection between the $k$-subsets of $\mathbb{N}_{n}$ such that no two elements of the subset differ by an element of $\mathcal{Q}=\{1,m\}$, where $m=2,3,\ldots$, and the permutations $\pi$ of $\mathbb{N}_{n+1}$ that have $k$ pairs of digits exchanged such that (a) the pairs do not overlap and (b) for all $i=1,\ldots,n+2-m$ the permutation satisfies $\max_{j=0,\ldots,m-1}\pi(i+j)-\min_{j=0,\ldots,m-1}\pi(i+j)\leq m$.
@@
The permutation $\pi$ of $\mathbb{N}_{n+1}$ corresponding to a subset $\mathcal{S}$ of $\mathbb{N}_{n}$ is formed as follows. If $i\in\mathcal{S}$ (where $i=1,\ldots,n$) then $\pi(i)=i+1$ and $\pi(i+1)=i$ (i.e., a pair of exchanged digits) whereas if neither $i$ nor $i-1$ are in $\mathcal{S}$ (for $i=2,\ldots,n+1$) then $\pi(i)=i$. As $\mathcal{Q}$ contains 1, the pairs of exchanged digits in the permutation cannot overlap. Given that condition (a) is satisfied, the only way for (b) to be violated would be if $\pi(i+1)=i$ and $\pi(i+m)=i+m+1$; this could only occur if both $i$ and $i+m$ were in $\mathcal{S}$ which is impossible since $m\in\mathcal{Q}$. The process of forming a permutation from a subset is clearly reversible and so the bijection is established. ∎
Theorem 8 has connections to the following sequences in the OEIS [20]: A000930 ($S^{\{1,2\}}_{n}$) and A102547 ($S^{\{1,2\}}_{n,k}$), its associated triangle; A130137 ($S^{\{1,3\}}_{n}$), which we mention again in the section on bit strings; A263710 ($S^{\{1,4\}}_{n}$), which was the sequence that led to the discovery of the bijection and we will now prove the conjectured generating function for that entry; and A374737 ($S^{\{1,5\}}_{n}$), which was recently to the OEIS added by the author.
@@
Figure 3: Digraph for restricted-overlap tiling with squares ($S$) and $(2,2,1)$-combs ($C$), which corresponds to the $\mathcal{Q}=\{1,4\}$ case.
@@
$g^{\{1,4\}}(x)=\frac{1}{x^{4}}\biggl{(}\frac{1-x^{3}-x^{5}}{1-x^{3}-x^{5}-x+x^{4}+x^{6}-x^{5}-x^{7}}-\frac{1-x^{4}}{1-x}\biggr{)},$
@@
$G(x)=\frac{1}{x^{5}}\left(\frac{\displaystyle{1}}{\displaystyle{1-x-\frac{\displaystyle{\frac{\displaystyle{x^{6}}}{\displaystyle{1-x^{3}}}+\frac{\displaystyle{x^{8}}}{\displaystyle{1-x^{2}}}}}{\displaystyle{1-\frac{\displaystyle{x^{4}}}{\displaystyle{1-x^{3}}}-\frac{\displaystyle{x^{6}}}{\displaystyle{1-x^{2}}}}}}}-\frac{1-x^{5}}{1-x}\right),$
@@
The proof we give of the second bijection (Theorem 12) requires the following bijection established in [9] between strongly restricted permutations and tiling with $(\frac{1}{2},g)$-fences (which are just $(\frac{1}{2},g,\frac{1}{2})$-combs) where $g$ is always a nonnegative integer. As the teeth (originally called posts in the context of fences) are not of integer width, it is convenient to regard each cell as being divided into two equal slots into which a tooth can fit [10].

### v3

@@ Abstract
Let $S_{n}$ and $S_{n,k}$ be, respectively, the number of subsets and $k$-subsets of $\mathbb{N}_{n}=\{1,\ldots,n\}$ such that no two subset elements differ by an element of the set $\mathcal{Q}$, the largest element of which is $q$. We prove a bijection between such $k$-subsets when $\mathcal{Q}=\{m,2m,\ldots,jm\}$ with $j,m>0$ and permutations $\pi$ of $\mathbb{N}_{n+jm}$ with $k$ excedances satisfying $\pi(i)-i\in\{-m,0,jm\}$ for all $i\in\mathbb{N}_{n+jm}$. We also identify a bijection between another class of restricted permutation and the cases $\mathcal{Q}=\{1,q\}$ and obtain generating functions for $S_{n}$ when $q=4,5,6$. We give some classes of $\mathcal{Q}$ for which $S_{n}$ is also the number of compositions of $n+q$ into a given set of allowed parts. A bijection between the $k$-subsets for any $\mathcal{Q}$ and bit strings is also noted. Aided by this, an efficient algorithm for finding $S_{n}$ and $S_{n,k}$ is given. We also prove a bijection between $k$-subsets for a class of $\mathcal{Q}$ and the set representations of size $k$ of equivalence classes for the occurrence of a given length-($q+1$) subword within bit strings. We then formulate a straightforward procedure for obtaining the generating function for the number of such equivalence classes.
@@ 1 Introduction
When there are no restrictions, we just have the number of ordinary combinations and so $S^{\{\}}_{n}=2^{n}$ (A000079 in the OEIS [20]) and $S^{\{\}}_{n,k}=\tbinom{n}{k}$ (A007318). It appears to be well-known that $S^{\{1,\ldots,q\}}_{n}=S^{\{1,\ldots,q\}}_{n-1}+S^{\{1,\ldots,q\}}_{n-q-1}+\delta_{n+q,0}$, where $\delta_{i,j}$ is 1 if $i=j$ and 0 otherwise, and $S^{\{1,\ldots,q\}}_{n,k}=\tbinom{n+q(1-k)}{k}$ (A329146). These reduce to the classic results $S^{\{1\}}_{n}=F_{n+2}$, where $F_{n}$ is the $n$-th Fibonacci number (A000045, $F_{n}=F_{n-1}+F_{n-2}+\delta_{n,1}$, $F_{n<1}=0$), and $S^{\{1\}}_{n,k}=\tbinom{n+1-k}{k}$ (A011973) [10]. Expressions for $S_{n}^{\mathcal{Q}}$ and $S^{\mathcal{Q}}_{n,k}$ have also been obtained when $\mathcal{Q}=\{q\}$ [13, 19, 15, 14, 3] and $\mathcal{Q}=\{m,2m,\ldots,jm\}$ with $j,m>0$ [17, 1].
Recently, a bijection between the $k$-subsets of $\mathbb{N}_{n}$ with disallowed differences set $\mathcal{Q}$ of which the largest element is $q$ and the restricted-overlap tilings of an $(n+q)$-board (an $(n+q)\times 1$ board) with squares and $k$ combs was found. This enabled generating functions for $S^{\mathcal{Q}}_{n}$ and recursion relations for $S^{\mathcal{Q}}_{n,k}$ to be readily obtained for a variety of families of $\mathcal{Q}$ [2]. A $(w_{1},g_{1},w_{2},\ldots,g_{t-1},w_{t})$-comb is a tile composed of $t>0$ rectangular sub-tiles (called teeth) arranged in such a way that the $i$-th tooth, which has dimensions $w_{i}\times 1$, is separated from the $(i+1)$-th by a $g_{i}\times 1$ gap [2]. The length of a comb is the sum of the lengths of the teeth and gaps. Each $\mathcal{Q}$ corresponds to a length-$(q+1)$ comb $C$ (which we will refer to as a $\mathcal{Q}$-comb) defined as follows. We label the cells of $C$ from 0 to $q$. Cell 0 is always part of the first tooth. Cell $i$ (where $i=1,\ldots,q$) is part of a tooth iff $i\in\mathcal{Q}$. Note that we can always include 0 in $\mathcal{Q}$ without affecting the results (and we will do so in Theorems 16 and 17 for reasons of convenience and symmetry) since subset elements are distinct and so cannot differ by 0. This 0 element can then be regarded as corresponding to cell 0 of the comb. To ensure that there is only one way to specify the $w_{i}$ and $g_{i}$ of the $\mathcal{Q}$-comb, we insist that all teeth and gaps have positive integer length. A restricted-overlap tiling using squares and $\mathcal{Q}$-combs means that any comb can be placed so that its non-leftmost cells (i.e, cells 1 to $q$) overlap the non-leftmost cells of any other combs on the board and no other forms of overlap of tiles is allowed [2] (see Fig. 7 for examples). We let $B^{\mathcal{Q}}_{n}$ denote the number of restricted-overlap tilings of an $n$-board with $\mathcal{Q}$-combs and squares, and $B^{\mathcal{Q}}_{n,k}$ denote the number of such tilings that contain $k$ combs. The following theorem is a consequence of the bijection.
@@
For all nonnegative $n$ we have $S^{\mathcal{Q}}_{n}=B^{\mathcal{Q}}_{n+q}$ and $S^{\mathcal{Q}}_{n,k}=B^{\mathcal{Q}}_{n+q,k}$.
@@
A cycle that does not include the 0 node is called an inner cycle (e.g., $SSC_{[4]}$ in Fig. 4, where including the length increment subscript quickly identifies which comb arc we are referring to). There are a finite number of possible metatiles iff the digraph lacks an inner cycle since a walk can traverse an inner cycle an arbitrary number of times before returning to the 0 node. If all the inner cycles have at least one node in common, that node is called a common node (e.g., the $01^{q-p}$ node is the common node in the Fig. 8 digraph; either the $10^{2}$ 1 or the 01 node can be chosen as the common node in the digraph in Fig. 3; Fig. 4 and Fig. 5 show digraphs that have inner cycles but no common node). For digraphs with a common node, there are simple expressions for the recursion relations for the numbers of tilings $B^{\mathcal{Q}}_{n}$ and $B^{\mathcal{Q}}_{n,k}$ in terms of lengths of various cycles in the digraph [8, 2] (we rederive these in the form of generating functions in §2). If the digraph possesses inner cycles but no common node, there is a class of digraph (which has 3 inner cycles) where an analogous simple procedure has been found for obtaining the recursion relations [1].
@@
A systematic search of the OEIS for sequences $(S^{\mathcal{Q}}_{n})_{n\geq 0}$ for all $\mathcal{Q}$ with $q\leq 4$ showed that there are various connections between some types of restricted combinations and some types of restricted permutations. Relations to compositions and bit strings were also found. Here, in §3, we give bijections between between two classes of strongly restricted permutations and combinations with $\mathcal{Q}=\{1,q\}$ and $\mathcal{Q}=\{m,2m,\ldots,jm\}$. As a consequence of the former bijection, we can prove a conjectured recursion relation for the permutations corresponding to the $\mathcal{Q}=\{1,4\}$ case. In §2, we derive generating functions for $B^{\mathcal{Q}}_{n}$ and $B^{\mathcal{Q}}_{n,k}$ for tilings with metatiles generated by a particular class of digraph (which includes the 3 inner cycle class with no common node dealt with previously [1]) without a common node. This enables us to also obtain a generating function for $S^{\{1,5\}}_{n}$. In §4, although it is not all that surprising given the tiling interpretation of restricted combinations, for completeness, we also show that $S^{\mathcal{Q}}_{n}$ is the number of compositions of $n+q$ into certain sets of parts for some classes of $\mathcal{Q}$. An elementary connection between all instances of $\mathcal{Q}$ and bit strings is pointed out in §5. This forms the basis for an algorithm for efficiently calculating $S^{\mathcal{Q}}_{n}$ and $S^{\mathcal{Q}}_{n,k}$ for any given $\mathcal{Q}$ as detailed via a heavily annotated C program listed in the Appendix. In §5 we also give a bijection between $k$-subsets for certain $\mathcal{Q}$ and equivalence classes for the occurrence of certain subwords within bit strings. The bijection enables us to show that a generating function for the number of equivalence classes is always straightforward to obtain.
@@ 2 Generating functions for numbers of walks on digraphs with a pseudo-common node
A form of the following theorem was originally stated in [8], and a more compact version re-derived in [2]. Here, using the above three lemmas, we obtain the result in terms of a bivariate generating function. We will use this theorem in §3.1 to obtain the $S^{\{1,4\}}_{n}$ generating function.
@@
To obtain the generating functions for $S^{\mathcal{Q}}_{n}$ and $S^{\mathcal{Q}}_{n,k}$, which we denote by $g^{\mathcal{Q}}(x)$ and $g^{\mathcal{Q}}(x,y)$, respectively, we require the following result.
@@
If $G(x,y)$ is the bivariate generating function for $B^{\mathcal{Q}}_{n,k}$ then
@@ 3 Bijections between strongly restricted permutations and restricted combinations
A strongly restricted permutation $\pi$ of the set $\mathbb{N}_{n}$ is a permutation for which the number of permissible values of $\pi(i)-i$ for each $i\in\mathbb{N}_{n}$ is less than a finite number independent of $n$ [16]. Here we give two bijections between restricted combinations and types of strongly restricted permutations where the permutations satisfy $\pi(i)-i\in\mathcal{D}$.
3.1 Connection with the class $\mathcal{Q}=\{1,q\}$
Our first bijection concerns a strongly restricted permutation where $\mathcal{D}=\{-1,0,1\}$ and no two of $q$ consecutive $\pi(i)$ may differ from one another by more than $q$.
@@ 3.1 Connection with the class $\mathcal{Q}=\{1,q\}$
There is a bijection between the $k$-subsets of $\mathbb{N}_{n}$ such that no two elements of the subset differ by an element of $\mathcal{Q}=\{1,q\}$, where $q=2,3,\ldots$, and the permutations $\pi$ of $\mathbb{N}_{n+1}$ that have $k$ pairs of consecutive items that have between exchanged such that for all $i=1,\ldots,n+2-q$ the condition $\max_{j=0,\ldots,q-1}\pi(i+j)-\min_{j=0,\ldots,q-1}\pi(i+j)\leq q$ is met.
@@
The permutation $\pi$ of $\mathbb{N}_{n+1}$ corresponding to an allowed subset $\mathcal{S}$ of $\mathbb{N}_{n}$ where $\mathcal{Q}=\{1,q\}$ is formed as follows. If $i\in\mathcal{S}$ (where $i=1,\ldots,n$) then $\pi(i)=i+1$ and $\pi(i+1)=i$ (i.e., a pair of two consecutive items has been exchanged) whereas if neither $i$ nor $i-1$ are in $\mathcal{S}$ (for $i=2,\ldots,n+1$) then $\pi(i)=i$. As $\mathcal{Q}$ contains 1, once a pair of items has been exchanged, neither item will be moved again. Given that $\pi(i)$ can be no more than 1 away from $i$, the only way that the condition could be violated would be if $\pi(i+1)=i$ (and hence $\pi(i)=i+1$) and $\pi(i+q)=i+q+1$; this could only occur if both $i$ and $i+q$ were in $\mathcal{S}$ which is impossible since $q\in\mathcal{Q}$. The process of forming a permutation from a subset is clearly reversible and so the bijection is established. ∎
For example, the allowed subsets of $\mathbb{N}_{5}$ when $\mathcal{Q}=\{1,3\}$ followed by the corresponding permutation of $\mathbb{N}_{6}$ are {} 123456, {1} 213456, {2} 132456, {3} 124356, {4} 123546, {5} 123465, {1,3} 214356, {2,4} 132546, {3,5} 124365, {1,5} 213465, and {1,3,5} 214365.
Theorem 8 has connections to the following sequences in the OEIS [20]: A000930 ($S^{\{1,2\}}_{n}$) and A102547 ($S^{\{1,2\}}_{n,k}$), its associated triangle; A130137 ($S^{\{1,3\}}_{n}$), which we mention again in the section on bit strings; A263710 ($S^{\{1,4\}}_{n}$), which was the sequence that led to the discovery of the bijection and we will now prove the conjectured generating function for that entry; along with A374737 ($S^{\{1,5\}}_{n}$) and A385870 ($S^{\{1,6\}}_{n}$), which were recently added to the OEIS by the author.
@@
Figure 3: Digraph for restricted-overlap tiling with squares ($S$) and $(2,2,1)$-combs ($C$), which corresponds to the $\mathcal{Q}=\{1,4\}$ case. In this and subsequent figures showing digraphs, each number in square brackets gives the contribution to the length of the metatile resulting from the addition of the tile. This subscript is omitted if there is no contribution, as is the case for any square placed in a gap of a comb.
@@
$g^{\{1,4\}}(x)=\frac{1}{x^{4}}\biggl{(}\frac{1-x^{3}-x^{5}}{1-x^{3}-x^{5}-x+x^{4}+x^{6}-x^{5}-x^{7}}-\frac{1-x^{4}}{1-x}\biggr{)},$ (4)
@@
$g^{\{1,5\}}(x)=\frac{1}{x^{5}}\left(\frac{\displaystyle{1}}{\displaystyle{1-x-\frac{\displaystyle{\frac{\displaystyle{x^{6}}}{\displaystyle{1-x^{3}}}+\frac{\displaystyle{x^{8}}}{\displaystyle{1-x^{2}}}}}{\displaystyle{1-\frac{\displaystyle{x^{4}}}{\displaystyle{1-x^{3}}}-\frac{\displaystyle{x^{6}}}{\displaystyle{1-x^{2}}}}}}}-\frac{1-x^{5}}{1-x}\right),$
@@
The metatile-generating digraph corresponding to the next case in the $\{1,q\}$ series (Fig. 5) does not fall into any of the classes of digraph that we have derived a generating function for so far. However, as we will now show, the systematic application of the digraph generating function rules (Lemmas 2, 3, and 4) can still be used to yield the generating function in this instance. To assist in breaking the derivation into more manageable chunks, we introduce the following notation. We let $\operatorname{\mathit{G}_{X\to Y}^{\{Z\}}}$ denote the generating function corresponding to all possible walks from node $X$ to node $Y$ that do not include node $X$ a second time and never encounter node $Z$. We let $\operatorname{\mathit{G}_{X}^{\{Z\}}}$ denote the generating function corresponding to all possible walks from node $X$ to itself without visiting itself or node $Z$ in between.
Proposition 11.
The generating function $g^{\{1,6\}}(x)$ for the number of restricted combinations of $\mathbb{N}_{n}$ with disallowed differences $\{1,6\}$ is given by
[…]
Figure 5: Digraph for restricted-overlap tiling with squares ($S$) and $(2,4,1)$-combs ($C$), which corresponds to the $\mathcal{Q}=\{1,6\}$ case. A comma-separated list of labels at an arc signifies a multiarc with those labels.
Proof.
The generating function $G(x,y)$ corresponding to paths starting at the 0 node of the digraph in Figure 5 satisfies
$\frac{1}{G(x,y)}=1-x-\frac{x^{7}y\operatorname{\mathit{G}_{0^{4}1\to 01}^{\{0\}}}}{1-\operatorname{\mathit{G}_{0^{4}1}^{\{0\}}}},$
where, owing to the fact that the only way to return to the $10^{4}$ 1 node is via the $C_{[5]}$ arc from the 01 node, we have $\operatorname{\mathit{G}_{0^{4}1}^{\{0\}}}=x^{5}y\operatorname{\mathit{G}_{0^{4}1\to 01}^{\{0\}}}$. The arcs leaving the $10^{4}$ 1 node are $S$ and $C_{[2]}$ (the latter contributes a factor of $x^{2}y$ to the corresponding term in the generating function) and the only way to reach the 01 node is from the $10^{2}$ 1 and 0101 nodes. Hence
$\operatorname{\mathit{G}_{0^{4}1\to 01}^{\{0\}}}=\frac{\operatorname{\mathit{G}_{0^{3}1\to 0^{2}1}^{\{01\}}}+\operatorname{\mathit{G}_{0^{3}1\to 0101}^{\{01\}}}}{1-\operatorname{\mathit{G}_{0^{3}1}^{\{01\}}}}+\frac{x^{2}y\bigl{(}\operatorname{\mathit{G}_{0^{2}101\to 0^{2}1}^{\{01\}}}+\operatorname{\mathit{G}_{0^{2}101\to 0101}^{\{01\}}}\bigr{)}}{1-\operatorname{\mathit{G}_{0^{2}101}^{\{01\}}}}.$
There is a direct route from $10^{3}$ 1 to $10^{2}$ 1 along with those that leave $10^{3}$ 1 via the $C_{[3]}$ arc (which contributes a factor of $x^{3}y$). On reaching $1010^{2}$ 1 there are two possible loops it can execute ($C_{[2]}SC_{[3]}$ or $C_{[2]}^{2}C_{[3]}$) before reaching $10^{2}$ 1. Hence
$\operatorname{\mathit{G}_{0^{3}1\to 0^{2}1}^{\{01\}}}=1+\frac{x^{3}y}{1-x^{5}y^{2}-x^{7}y^{3}}.$
Since the only allowed route returning to the $10^{3}$ 1 node is the $C_{[4]}$ arc from the $10^{2}$ 1 node, $\operatorname{\mathit{G}_{0^{3}1}^{\{01\}}}=x^{4}y\operatorname{\mathit{G}_{0^{3}1\to 0^{2}1}^{\{01\}}}$. In a similar fashion we obtain
$\operatorname{\mathit{G}_{0^{3}1\to 0101}^{\{01\}}}=\frac{x^{5}y^{2}(1+x^{2}y)}{1-x^{5}y^{2}-x^{7}y^{3}},\qquad\operatorname{\mathit{G}_{0^{2}101\to 0^{2}1}^{\{01\}}}=\frac{(1+x^{2}y)x^{3}y}{1-x^{4}y-x^{7}y^{2}},$
and $\operatorname{\mathit{G}_{0^{2}101\to 0^{2}1}^{\{01\}}}=1+x^{2}y$. The $1-x^{7}y^{2}/(1-x^{4}y)$ in
$\operatorname{\mathit{G}_{0^{2}101}^{\{01\}}}=\frac{\displaystyle{(1+x^{2}y)x^{5}y^{2}}}{\displaystyle{1-\frac{x^{7}y^{2}}{1-x^{4}y}}}$
results from a loop within a loop: the $SC_{[4]}$ loop can be executed on arriving at the $10^{3}$ 1 node while traversing the $SC_{[4]}C_{[3]}$ loop. Combining the above results to give $G(x,y)$ and applying Corollary 7 to $G(x,1)$ yields (6) after simplifying. ∎
3.2 Connection with the class $\mathcal{Q}=\{m,2m,\ldots,jm\}$
The proof we give of the second bijection (Theorem 13) requires the following bijection established in [8] between strongly restricted permutations and tiling with $(\frac{1}{2},g)$-fences (which are just $(\frac{1}{2},g,\frac{1}{2})$-combs) where $g$ is always a nonnegative integer. As the teeth (originally called posts in the context of fences) are not of integer width, it is convenient to regard each cell as being divided into two slots into which a tooth can fit [9].